OpenStudy (korosh23):

Pre-calculus 12 Question! Trigonometry question! radius or r= 10 Angle and coordinates are in in quadrant 2 sin theta= 3/5 Determine the coordinates.

OpenStudy (korosh23):

sin theta= y/r y= 3 or 6 r= 5 or 10 x= by pythagreom theorem

OpenStudy (korosh23):

x= 8 I am wondering if the coordinates are (-4/5 , 3/5) or ( -8, 6)

OpenStudy (johnweldon1993):

Remember the coordinates \(\large (rcos(\theta), rsin(\theta))\)

OpenStudy (korosh23):



OpenStudy (korosh23):

Oh interesting, so that means the coordinates are (-8 , 6)

OpenStudy (johnweldon1993):

Correct indeed :)

OpenStudy (korosh23):

Thank you :)

OpenStudy (johnweldon1993):

No problem :)